A 1500 g object is being pushed by an applied force and accelerates at 4.08 m/s2. what is the force applied if a frictional force is 62.5 n is acting on opposite direction?


1
Expert's answer
2021-11-07T19:26:25-0500

By Newton's second law:


"ma=F-f,\\
F=ma+f=1.5·4.08+62.5=68.62\text{ N}."